# Can industrial overcapacity enable seasonal flexibility in electricity use? A case study of aluminum smelting in China

### v1

@@ 2.3 Cost-benefit analysis of retaining overcapacity
Figure 4: Trade-off analysis and cost-effective strategy for retaining overcapacity. In (a), electricity system cost savings (blue), smelter operational cost increase (orange), and net system benefit (black line) are shown for 2050 in the core scenario (Mid-flexibility, Mid-demand, and Mid-technology costs). The maximal net benefit (electricity system cost reduction minus smelter operational costs) is achieved with 36% overcapacity in 2050. In (b), the mean optimal overcapacity rate increases from 6% in 2030 to 30% in 2050, due to higher flexibility requirements as the electricity system approaches net-zero emissions. However, it should be noted that the cost-effective capacity of aluminum smelting still decreases across the years due to declining primary aluminum demand. The net benefit across different scenarios first decreases as primary aluminum demand declines to 2040 levels, then increases in the net-zero emissions scenario in 2050, where load flexibility is more valuable.
While flexible aluminum smelting with overcapacity can reduce power system costs, it also entails higher maintenance costs and extra expenditure, such as product storage. As shown in Figure 4(a), the reduced power system costs increase with overcapacity but plateau after smelter capacity exceeds 35 Mt/year (60% overcapacity), while operational costs increase approximately proportionally with excess capacity. A smelting capacity of 18 Mt/year (36% overcapacity) is found to maximize the net system cost reduction at 17 billion CNY, where the marginal decrease in power system cost equals the marginal increase in smelter operational cost.
Figure 4(a) also shows the impact of smelter load flexibility for carbon emission reduction, indicated by the red line. The reduction in carbon emissions represents a decrease in the system’s need for negative emission technologies, such as biomass with carbon capture and storage (CCS). This value continues to increase with overcapacity when it exceeds 36%, implying that more maintenance costs can be spent to achieve carbon reduction benefits. Retaining more excess capacity could be an even more attractive option if the cost of retiring capacity is taken into account (dismantling retired electrolytic cells, compensating affected smelters, and relocating displaced industrial workers, etc.). These additional costs could further justify maintaining higher levels of excess capacity beyond what our results suggest.
Figure 4(b) shows the optimal smelter capacity and corresponding net system cost reduction for different years and scenarios. The average optimal overcapacity rates across all scenarios are 6%, 20%, and 30% for 2030, 2040, and 2050, respectively. The corresponding net value first decreases from 2030 to 2040, as the primary aluminum industry demand declines to 2040 levels, then increases by 2050, approaching net-zero emission targets.
@@ 2.4 Seasonal operation paradigm enabled by overcapacity
(a) Normalized hourly aluminum smelting power consumption and stored aluminum with 36% overcapacity in 2050.
(b) Monthly capacity factors of various load and generation resources in 2050. (Capacity Factor = energy generation/consumption of a technology, divided by its nameplate capacity)
@@
Figure 5: Seasonal operation of aluminum smelters is complementary to energy system seasonality. In (a), smelter production largely ceases from mid-November to mid-March with demand met by stored inventory, but operates at full capacity from late March to early November. In (b), the electricity consumption of aluminum smelting is largely complementary to China’s heating demand, avoiding the severe winter electricity peaks driven by heating electrification.
@@
Figure 5 illustrates how the seasonal nature of renewable energy generation and heating electrification significantly alters the operational dynamics of aluminum smelters, which adjust their production schedules to align with the seasonality of the energy system. Panel (a) reveals clear seasonal operational patterns: production largely ceases from mid-November to mid-March (the winter heating season), with demand met by inventory, while operating at or near rated capacity from late March to early November. The lower panel shows that even with unconstrained flexibility, seasonal operational patterns remain dominant in aluminum smelting, indicating that system needs for seasonal flexibility are the primary driver. Although China’s energy system faces significant cooling demand in July and August, smelters maintain maximum power consumption because air conditioning load appears as short-duration peaks that can be balanced by batteries or vehicle-to-grid (V2G) technologies, making smelter load reduction less valuable.
Panel (b) of Figure 5 reveals the underlying reason for flexible aluminum smelting: smelter electricity consumption is largely complementary to China’s overall heating demand, strategically avoiding high winter electricity demand caused by heating electrification. During these winter periods, renewable energy output, especially from hydropower, is at its minimum while heating demand peaks, forcing the power system to rely more heavily on gas and coal generation to ensure power supply reliability. This results in significantly higher electricity costs, because the system must deploy more direct air capture and/or bioenergy with CCS to offset the carbon emissions, creating strong economic incentives for smelters to avoid production during these high-cost periods and instead rely on inventory to meet demand. Interestingly, the averaged optimal overcapacity rate of 30% roughly corresponds to the proportion of winter time in a year, suggesting that seasonality is the driver behind this.
@@
Figure 6: Comparison of the composition of levelized cost per tonne of aluminum between the $20$ baseline and various 2050 overcapacity levels. The composition of the levelized cost per tonne of aluminum is shown across three overcapacity levels in the core scenario (Mid-flexibility, Mid-demand, and Mid-technology costs) and compared against a 2020 baseline. We find that maintaining 36% overcapacity is beneficial for the smelters, reducing aluminum production costs by over 1,500 CNY/tonne, or approximately 9% of current production costs, with the drop in electricity bills offsetting the maintenance costs of overcapacity and increased restart and storage costs.
@@
This seasonal operational paradigm could decrease the cost of producing aluminum (Figure 6). If the historical overcapacity rate of 5% is retained, the average cost per tonne of aluminum may slightly increase by 2050. Although electricity costs decrease due to increased renewable penetration, total costs remain the same due to equipment depreciation costs and early decommissioning costs. In contrast, maintaining 36% overcapacity allows smelters to consume zero-marginal-cost renewable energy during high-generation periods, thus largely reducing electricity bills to offset overcapacity maintenance and product storage costs, and reducing total costs by over 1,500 CNY/tonne of aluminum. The reduced aluminum production costs, along with the electricity system cost savings shown previously, demonstrate that the seasonal flexibility strategy creates a win-win scenario for both the grid and aluminum producers, altering the economic viability of maintaining excess capacity.
Figure 7: Estimated monthly workforce requirements for aluminum smelters and coal/gas power plants in 2050 with and without aluminum smelting overcapacity. In the overcapacity scenario, complementary employment patterns between smelters and thermal generators reduce workforce variability (calculated as the standard deviation of employment throughout the year) by 62% and increase average employment by 18,000, as compared to the scenario in which all excess capacity is retired.
@@
As Figure 7 illustrates, the aluminum smelting industry, with overcapacity, and the power sector have complementary workforce patterns, with coal and gas power plants needing more workers in the winter, and the aluminum smelting industry needing more workers in the summer. Because China’s largest aluminum companies have historically operated their own captive power plants, there exists a possible scenario in which China’s aluminum companies move workers between aluminum smelters and power plants from season to season [tan_different_2025]. This workforce management paradigm could reduce total seasonal employment fluctuation across the two sectors by 62%, while maintaining a larger capacity also increases average employment by 18,000 workers compared to the no-overcapacity scenario. While implementing inter-industry labor mobility depends on factors such as worker training and flexible employment relations, this coupling could potentially mitigate social disruption from job losses that may result from both energy system transitions and industrial restructuring, highlighting the potential societal value of overcapacity-enabled flexibility.
@@ 3 Discussion
In power systems with increasing penetration of renewable energy, the volatility of electricity supply is often compounded by load fluctuations from heating electrification, creating significant seasonal variations in electricity costs. Concurrently, many traditional energy-intensive industries face structural overcapacity, which reduces capacity utilization and creates uncertainty on their profitability. This convergence of challenges makes it increasingly urgent to understand how energy-intensive industries can adapt and thrive in decarbonized energy systems.
@@
In this work, we use China’s aluminum smelting industry—a highly energy-intensive sector facing unprecedented overcapacity in the coming recycling era—as a representative case to explore the flexibility value of overcapacity in decarbonized energy systems. We developed a comprehensive framework that co-optimizes power system capacity expansion planning with detailed smelter operations, capturing key technical constraints such as temperature limits, restart costs, and operational limitations. This co-optimization was evaluated across more than 500 scenarios, varying flexibility settings and competing technology costs, to inform cost-effective strategies for retaining overcapacity and leveraging it for seasonal flexibility.
@@
Our findings reveal that overcapacity unlocks seasonal flexibility of smelting electricity use, allowing them to curtail production during the winter when peak demand is exacerbated by heating electrification and reduced renewable output. In the 2050-net-zero scenario, seasonal flexibility reduces annual power system costs by 15-72 billion CNY through improved renewable integration and decreased reliance on firm generation, sufficient to offset the maintenance and operational costs of excess capacity and product storage, thereby cutting aluminum production costs by approximately 9% of the current level. We identify an optimal overcapacity rate of approximately 30% for the aluminum smelting industry in China, which closely matches the fraction of the year that is the winter heating season, indicating that the seasonality of the energy system is the fundamental driver of a seasonal smelter operation paradigm.
Our findings suggest that the increasing seasonality of decarbonized energy systems can be strong enough to drive energy-intensive industries to shift to a seasonal operation paradigm. Energy-intensive industries, such as aluminum, steel, and cement, will be the first to experience these impacts, given their high exposure to energy cost variations and structural overcapacity. As electricity cost variability becomes more seasonal, these sectors can leverage overcapacity and product storage capabilities to adapt to volatile electricity prices, reducing both their own operational expenses and overall energy system costs.
Interestingly, overcapacity can facilitate complementary workforce patterns between energy-intensive industries and the energy sector, which can help mitigate social disruption from energy decarbonization and industrial restructuring. Nevertheless, converting the co-benefits of overcapacity into financial incentives for energy-intensive industries will require careful design of financial instruments, targeted subsidies or market mechanisms, as is required for the seasonally operating agriculture sector.
@@ 4 Methods
4.1 Scenario analysis framework
@@
The factors that have significant and uncertain impacts on the value of overcapacity in the aluminum industry include the demand for primary aluminum (from smelting), the operational flexibility of the smelters, and the costs of competitive technologies for load flexibility (batteries, long-duration storage, etc.). This study employs a scenario analysis approach, grouping these factors into three-dimensional parameter sets (Table 1) to explore the impact of different assumptions and uncertain parameters on the research findings.
In the model, China’s domestic aluminum smelters are required to meet the country’s annual primary aluminum demand, which is equal to the total aluminum demand minus the demand that can be met by recycled aluminum, as recycled aluminum is prioritized for production due to its per-unit mineral, energy, and investment costs being orders of magnitude lower than those for aluminum smelting. We focus exclusively on primary aluminum smelters and do not consider recycled aluminum production to be overcapacity, as its output is currently increasing and does not face overcapacity issues. Additionally, recycled aluminum production has limited grid interaction flexibility value due to its low electricity intensity (only 5% of the electricity used in primary aluminum production) and is not traditionally considered an energy-intensive industry.
@@
Using performance data from numerical simulations of aluminum smelter operations, as well as existing literature on the shutdown, restart processes, and costs of real-world aluminum smelters, we developed a reduced-order model that represents the key characteristics of aluminum electrolytic cell temperature constraints and restart costs (see Supplementary note 3 for more details). However, the cost uncertainty is large because there are few historical cases of aluminum smelters being restarted. Thus, we evaluated three comparative scenarios: Low: No progress has been made in improving smelter operational flexibility, and restart costs from historical high-cost reports are being used. High: New technologies and methods have significantly improved smelter operational flexibility, and restart costs have dropped substantially. This represents the goal of ongoing research, development, and demonstration efforts to enhance smelter-grid interaction [lukin_reduction_2016]. Un-constrained: This is an idealized scenario that disregards the temperature constraints, startup/shutdown constraints, and restart costs of potlines. It only considers rated power constraints, and the value of grid interaction in this scenario represents the theoretical maximum.
@@
Finally, we consider three main power system technology cost scenarios (see Table 1) that reflect different levels of power generation technology development. All scenarios assume that a variety of competitive, clean, and stable technologies are available, covering a range from high fixed costs and low variable costs to low fixed costs and high variable costs. We also assume the availability of cost-competitive long-duration storage technologies (in the form of hydrogen and hot water) and a certain degree of flexibility in residential heating. By modeling combinations of different scenarios for primary aluminum demand, smelter flexibility, and technology costs, we are able to evaluate the sensitivity of the main power system and aluminum industry outcomes to changes along these dimensions.

### v2

@@ 2.3 Cost-benefit analysis of retaining smelting overcapacity
Figure 4: Trade-off analysis and cost-effective strategy for retaining overcapacity. In (a), electricity system cost savings, smelter cost increase, and net system benefit with different retained capacities in 2050 in the core scenario are shown. The maximal net benefit (electricity system cost reduction minus smelter increased costs) is achieved with 30% overcapacity in 2050. In (b), the mean optimal overcapacity rate increases from 6% in 2030 to 30% in 2050, due to higher flexibility requirements as the electricity system approaches net-zero emissions. However, it should be noted that the cost-effective overcapacity of aluminum smelting still decreases across the years due to declining primary aluminum demand. The net benefit across different scenarios first decreases as primary aluminum demand declines to 2040 levels, then increases in the net-zero emissions scenario in 2050, where load flexibility is more valuable.
While flexible aluminum smelting with overcapacity can reduce power system costs, it also entails higher maintenance costs, labor cost, restart cost, and extra expenditure, such as product storage. As shown in Figure 4(a), the reduced power system costs increase with overcapacity but plateau after smelter capacity exceeds 35 Mt/year (60% overcapacity), while operational costs increase approximately proportionally with excess capacity. A smelting capacity of 17 Mt/year (30% overcapacity) is found to maximize the net system cost reduction at 13 billion CNY, where the marginal decrease in power system cost equals the marginal increase in smelter operational cost. Retaining more excess capacity could be more attractive if the cost of retiring capacity is taken into account (dismantling retired electrolytic cells, compensating affected smelters, and relocating displaced industrial workers, etc.). These additional costs could further justify maintaining higher levels of excess capacity beyond what our results suggest.
Figure 4(b) shows the optimal smelter capacity and corresponding net system cost reduction for different years and scenarios. The average optimal overcapacity rates across all scenarios are 6%, 20%, and 30% for 2030, 2040, and 2050, respectively. The corresponding net value first decreases from 2030 to 2040, as the declined primary aluminum demand leads to a shrinkage of the aluminum smelting industry, then increases by 2050, approaching net-zero emission targets.
@@ 2.4 Seasonal operation paradigm enabled by overcapacity
(a) Normalized hourly aluminum smelting power consumption and stored aluminum with 30% overcapacity in 2050 core scenario.
(b) Monthly electricity demand and generation in 2050. (Net demand = electricity demand - electricity generation)
@@
Figure 5: Seasonal operation of aluminum smelters is complementary to energy system seasonality. In (a), smelter production largely ceases from mid-November to mid-March with demand met by stored inventory, but operates at full capacity from late March to early November. Comparing (a) with (b), the electricity consumption of aluminum smelting is largely complementary to China’s monthly net electricity demand, acting as a crucial seasonal flexibility resource to mitigate the winter electricity demand peaks driven by heating electrification.
@@
Figure 5 illustrates how the seasonal nature of renewable energy generation and heating electrification significantly alters the operational dynamics of aluminum smelters, which adjust their production schedules to align with the seasonality of the energy system. Panel (a) reveals clear seasonal operational patterns: production largely ceases from mid-November to mid-March (the winter heating season), with demand met by inventory, while operating at or near rated capacity from late March to early November. The lower panel shows that even with unconstrained aluminum smelter flexibility, seasonal operational patterns remain dominant in aluminum smelting. Although China faces significant cooling demand in July and August, smelters do not operate flexibly to balance that demand because air conditioning load appears as short-duration peaks that can more cost-effectively be balanced by batteries or vehicle-to-grid (V2G) technologies.
Panel (b) reveals the underlying system driver for this behavior: the emergence of a “seasonal duck curve”. Driven by high renewable penetration and heating electrification, China’s net electricity demand (residual demand after subtracting renewable generation) exhibits a distinct and prolonged winter peak. During these winter periods, renewable energy output, especially from hydropower, is at its minimum while heating demand peaks. Although wind power also increases in the winter, it is outpaced by the surge in heating demand, forcing the power system to rely more heavily on gas and coal generation to ensure power supply reliability. Here, the national aggregate solar output obscures regional differences (e.g., reduced summer solar generation in Yunnan due to the monsoon season). For provincial results, see Supplementary Figure 8.
@@
Figure 6: Comparison of the change in levelized cost per tonne of aluminum between the $20$ baseline and various 2050 overcapacity levels. The composition of the levelized cost per tonne of aluminum is shown across three overcapacity levels in the core scenario and compared against a 2020 baseline. We find that maintaining 30% overcapacity is beneficial for the smelters, reducing aluminum production costs by over 1,500 CNY/tonne (approximately 9% of current production costs), with the drop in electricity bills offsetting the maintenance costs of overcapacity and increased restart, storage, and capital costs.
This results in significantly higher electricity costs in winter, because the system must deploy more direct air capture and/or bioenergy with CCS to offset the carbon emissions. To balance this seasonal supply-demand mismatch and flatten the prolonged winter peak, the system requires dedicated seasonal flexibility resources. Aluminum smelting with overcapacity perfectly fills this role, creating strong economic incentives for smelters to avoid production during these high-cost periods and instead rely on inventory to meet demand. Interestingly, the averaged optimal overcapacity rate of 30% roughly corresponds to the proportion of winter time in a year, further confirming that this seasonal duck curve, rather than daily fluctuations, dictates the investment logic of overcapacity.
@@
This seasonal operational paradigm could decrease the cost of producing aluminum (Figure 6). Our analysis indicates that maintaining a moderate level of overcapacity (e.g., 30%, middle bar of Figure 6) minimizes overall production costs, outperforming both the complete decommissioning (leftmost bar of Figure 6) and full retention (rightmost bar of Figure 6) scenarios. By retaining 30% overcapacity, smelters can effectively utilize zero-marginal-cost renewable energy during high-generation periods. The resulting electricity cost savings are sufficient to offset the added expenses of overcapacity maintenance, incremental capital (see Supplementary Note 3.6), and product storage, ultimately reducing total costs by 9% (over 1,500 CNY/tonne of aluminum). In contrast, if no overcapacity is retained, the savings from general renewable penetration are largely negated by equipment depreciation and early decommissioning costs. Conversely, retaining 100% of the initial overcapacity introduces fixed maintenance and capital burdens that outweigh the benefits of seasonal flexibility in electricity use.
The reduced aluminum production costs, along with the electricity system cost savings shown previously, demonstrate that the seasonal flexibility strategy creates a win-win scenario for both the grid and aluminum producers, altering the economic viability of maintaining excess capacity. It is worth noting that our current model only captures savings from electricity procurement. As electricity market designs evolve to value system adequacy, industrial facilities operating seasonally will likely capture additional value through capacity markets. Consequently, the actual financial returns for aluminum smelters are expected to be even more optimistic than our current projections. The detailed cost composition underlying these figures is provided in Supplementary Note 3.4.
Figure 7: Estimated monthly workforce requirements for aluminum smelters and coal/gas power plants in 2050 with and without aluminum smelting overcapacity. The results are from the flexible labor scenario with Mid-primary aluminum demand, Mid-smelter flexibility, and Mid-power system technology cost. With 30% smelter overcapacity, there is a potential temporal complementarity of labor demand between smelters and thermal generators. If such potential is realized, complementary employment patterns between the two sectors could reduce workforce variability, further unleashing the social benefits of seasonal flexibility.
@@
As Figure 7 illustrates, the seasonally-operating aluminum smelting industry and the power sector have complementary workforce demand patterns, with coal and gas power plants needing more workers in the winter, and the aluminum smelting industry needing more workers in spring, summer, and fall. Because China’s largest aluminum companies have historically operated their own captive power plants (see Supplementary Note 3.8), there exists a possible scenario in which a subset of China’s aluminum companies move workers between aluminum smelters and power plants from season to season [53]. This workforce management paradigm could reduce total seasonal employment fluctuation (calculated as the standard deviation of monthly employment throughout the year) across the two sectors by 25%. Maintaining a larger capacity also increases overall employment compared to the no-overcapacity scenario. While implementing inter-industry labor mobility depends on factors such as worker training and flexible employment relations, this coupling could potentially mitigate social disruption from job losses that may result from both energy system transitions and industrial restructuring, highlighting the potential societal value of overcapacity-enabled seasonal flexibility.
@@ 3 Discussion
In power systems with increasing penetration of renewable energy, the volatility of electricity supply is often compounded by load fluctuations from heating electrification, creating significant seasonal variations in electricity supply and demand. Concurrently, many traditional energy-intensive industries face persistent overcapacity, which reduces capacity utilization and creates uncertainty on their profitability. This convergence of challenges makes it increasingly urgent to understand how energy-intensive industries can adapt and thrive in decarbonized energy systems.
@@
In this work, we use China’s aluminum smelting industry—a highly energy-intensive sector facing unprecedented overcapacity in the coming recycling era—as a representative case to explore the flexibility value of overcapacity in decarbonized energy systems. We developed a comprehensive framework that co-optimizes power system capacity expansion planning with detailed smelter operations, capturing key technical constraints such as temperature limits, restart costs, and operational limitations. This co-optimization was evaluated across more than 1000 scenarios, varying flexibility settings and competing technology costs, to inform cost-effective strategies for retaining overcapacity and leveraging it for seasonal flexibility.
@@
Our findings reveal that overcapacity unlocks the flexibility of aluminum smelting operation, allowing smelters to curtail production during the winter when peak demand is exacerbated by heating electrification and reduced renewable output. We identify an optimal overcapacity rate of approximately 30% for the aluminum smelting industry in China’s decarbonized electricity power system, which closely matches the fraction of the year that is the winter heating season, indicating that the seasonality of the energy system translates into seasonal electricity price signals, which ultimately drive the smelters’ seasonal operation. This seasonal operation transforms a flat industrial power demand into a seasonal load profile that reduces mismatch between electricity supply and demand. This avoids the need for massive investment and operational costs for expensive clean-firm power technologies, while simultaneously improving renewable energy integration. These reduced investment and operational costs translate into cheaper electricity costs, sufficient to offset the maintenance and operational costs of excess capacity and product storage, thereby cutting aluminum production costs by approximately 9% of the current level.
Crucially, our findings suggest a need to re-evaluate the traditional paradigm of demand response, which in existing literature predominantly targets short-term (hourly to daily) demand-side flexibility. We demonstrate that energy-intensive industrial sectors have strong potential to be advantageous for long-term, seasonal demand flexibility. Industrial processes like primary aluminum smelting are fundamentally designed for steady-state, continuous operation. Forcing these systems to provide high-frequency, short-term flexibility often incurs steep efficiency penalties, thermal fatigue of equipment, risks of cell shutdown, and high control costs. In contrast, leveraging overcapacity allows these facilities to operate continuously at optimal efficiency for most of the year, undergoing only planned, long-duration curtailments during extreme seasonal grid peaks. Therefore, when accounting for the strict requirements of stable industrial operations, this seasonal paradigm emerges as not only economically superior but also technically far more practical than conventional short-term demand response.
More broadly, our work suggests that industrial overcapacity should be reframed from a problem to an opportunity in a world where low-cost but variable renewable energy is becoming one of the cheapest and most abundant forms of primary energy available for industrial activity. Just as the global economy has reconfigured multiple times around the availability and characteristics of new low-cost primary energy sources, the ongoing shift to renewable electricity may entail new paradigms for industrial operations, plant design, and the geography of primary industries. While our work focuses on aluminum smelting as a case study, these findings are potentially generalizable to other electricity-intensive industries facing overcapacity, such as steel and cement. By embracing this new seasonally-operating paradigm, these sectors may also be able to leverage overcapacity and product storage to adapt to volatile seasonal electricity prices, reducing both their own operational expenses and overall energy system costs, while enabling more cost-effective mitigation of greenhouse gas emissions.
Realizing this seasonal paradigm also requires addressing its socio-economic implications, particularly concerning labor. However, because the labor intensity of modern primary aluminum production is exceptionally low (accounting for only 2%–3% of total costs [49]), the immense energy cost savings achieved during high-generation periods are more than sufficient to cover standby wages or training stipends during winter lulls. As financial incentives make the transition to seasonal operations economically compelling, exploring potential employment complementarity under this paradigm represents a promising avenue for future research and policy design. For instance, the winter production lull in aluminum smelting naturally coincides with the operational peaks of thermal power plants in a decarbonized grid. While facilitating such inter-industry job mobility requires specific spatial planning (e.g., large integrated industrial parks) and targeted institutional support, it is not without precedent; seasonal labor deployment has long been successfully managed in sectors like agriculture and tourism. Ultimately, supported by carefully designed market and policy mechanisms, the seasonal operation of energy-intensive industries holds the potential to unlock broader societal benefits, transforming structural overcapacity into a strategic asset for employment resilience and social stability.
@@ 4 Methods
4.1 Data Source and Scenario Settings
@@
The factors that have significant and uncertain impacts on the value of overcapacity in the aluminum industry include: the demand for primary aluminum (from smelting), the operational flexibility of the smelters, the costs of other low-carbon technologies for load flexibility (batteries, long-duration storage, etc.), and the flexibility of labor and employment. This study employs a scenario analysis approach, grouping these factors into four-dimensional parameter sets (Table 1) to explore the impact of different assumptions and uncertain parameters on the research findings.
In our model, China’s domestic aluminum smelters are required to meet the annual primary aluminum demand, defined as the total demand minus the availability of recycled aluminum. This primary demand is a critical factor determining the system-level value of flexibility and the optimal level of strategically retained overcapacity: intuitively, a higher primary demand implies a larger industrial base with greater cumulative potential for seasonal load shifting and peak shaving. We focus exclusively on primary aluminum smelters for seasonal flexibility, modeling recycled (secondary) aluminum production as a continuous baseload process. This constraint is driven by three techno-economic factors: (1) unlike the legacy overcapacity in primary smelting, the rapidly expanding recycling sector requires substantial new capital investment, necessitating high utilization rates to amortize marginal capital expenditures; (2) the low energy intensity of recycling (roughly 5% of primary production) yields electricity cost savings during peak months that are insufficient to offset the depreciation of idle capacity; and (3) the temporal rigidity of scrap supply and its high inventory holding costs render seasonal shutdowns economically unviable. By exploring various demand scenarios, we evaluate how the value of overcapacity evolves as the industry transitions from a primary-dominated era toward a recycling-dominated one.
@@
The flexibility value of aluminum smelting and its resulting operational patterns are inherently governed by the physical constraints of the smelting process, specifically the extent to which individual potlines can adjust their power consumption and the technical difficulty associated with complete shutdowns and restarts. Using performance data from numerical simulations of aluminum smelter operations, as well as existing literature on the shutdown, restart processes, and costs of real-world aluminum smelters, we developed a reduced-order model that represents the key characteristics of aluminum electrolytic cell temperature constraints and restart costs (see Supplementary Note 3 for more details). However, the cost uncertainty is large because there are few historical cases of aluminum smelters being restarted. Thus, we evaluated four comparative scenarios:
Low: Based on the traditional understanding of the aluminum smelting process, potlines must operate stably with almost no power adjustment capability except for a complete shutdown. Restart costs are based on historical high-cost reports.
Mid (Core Scenario): Based on recent observations in cases such as Yunnan, China, smelters can reduce the operating power of individual potlines by 10% without significant modifications. Restart costs are also derived from reported values in these Chinese cases.
High: New technologies and methods significantly improve smelter operational flexibility, allowing for power reductions of up to 30% and substantially lower restart costs. This represents the goal of ongoing research, development, and demonstration efforts to enhance smelter-grid interaction [36].
Un-constrained: This is an idealized scenario that disregards the temperature constraints, startup/shutdown constraints, and restart costs of potlines. It only considers rated power constraints, and the value of grid interaction in this scenario represents the theoretical maximum.
@@
The system-level flexibility value of aluminum smelting is also influenced by the costs of energy supply and the availability of alternative flexibility sources. Lower costs for renewable energy or competing flexibility technologies, such as batteries, could diminish the relative value of industrial load flexibility, thereby impacting the economic rationale for retaining overcapacity. To account for this, we consider three main power system technology cost scenarios (see Table 1) that reflect different levels of power generation technology development. All scenarios assume that a variety of competitive, clean, and stable technologies are available, covering a range from high fixed costs and low variable costs to low fixed costs and high variable costs. We also assume the availability of cost-competitive long-duration storage technologies (in the form of hydrogen and hot water) and a certain degree of flexibility in residential heating.
Finally, we introduce a scenario dimension concerning labor and employment flexibility, which may affect the economic feasibility of seasonal operation. We evaluate two comparative cases: (1) Inflexible labor (Core scenario): This represents the most unfavorable case where smelters must maintain their full workforce and associated labor and standby costs regardless of output levels or seasonal shutdowns. Workers are assumed to be unable to transfer to other industries. (2) Flexible labor: This case explores the potential for labor mobility and further cost optimization. If a potline is shut down, labor and standby costs are reduced. By modeling combinations of different scenarios for primary aluminum demand, smelter flexibility, technology costs, and labor flexibility, we are able to evaluate the sensitivity of the main power system and aluminum industry outcomes to changes along these dimensions.